# Incentivizing Reasoning for Advanced Instruction-Following of Large Language Models

### v1

@@ 4.1 Experimental Setup
Details on statistics about our self-evolving dataset can be found in Sec. A.3. Additionally, we incorporated the DeepScaleR deepscaler2025 maths dataset to facilitate reasoning development (see Sec. A.4). The evaluation metrics for each benchmark are reported in Sec. A.2.2.
@@
We compared with: 1) I/O: direct input with $x_{\text{CoT}}=\emptyset$; 2) CoT: reasoning prompting wei2022chain to first deliver the thought and then the answer; 3) SDC: self-DeepClaude ErlichLiu ; getAsterisk technique that first prompts for the thought and then packs the original input with the thought as a new context for the answer (see Sec. A.5.3); 4) SFT: supervised fine-tuning for learning the aligned responses.
@@
We use OpenRLHF hu2024openrlhf for both cold-start (Qwen2.5-1.5B/7B yang2024qwen2 , LLaMA3.1-8B grattafiori2024llama , and Ministral-8B jiang2023mistral7b ) and warm-start (DeepSeek-Qwen1.5B/7B guo2025deepseek and DeepScaleR-1.5B deepscaler2025 ) experiments. Detailed settings can be found in Sec. A.5. To ensure fairness and comparability, we employ Qwen2.5-72B-Instruct as the judge for all LLM-based evaluation.
@@ 4.2 Main Results
In line with Fig. 3, small models (1.5B) achieve much more gains than larger ones, showcasing the potentials of small LLMs via test-time scaling. The DeepSeek-distilled LLMs possess a good starting point for reasoning organization from their warm-start imitation across a broad range of tasks and topics (Fig. 36). With respect to model family, we unfortunately find that the capacity of Ministral and LLaMA is inferior to that of Qwen. The Ministral-8B exhibits limited advantages over its vanilla counterpart while the LLaMA3.1-8B experienced a model collapse during training. As shown in Fig. 35(a) and (d), a rapid shrinkage of response and a frequent surge of KL penalty imply a great deviation of LLaMA from its initial state. The reason behind might be ascribed to the pre-trained knowledge of base models gandhi2025cognitive . LLaMA tends to generate endless thinking without conforming to the required format. It then struggles to output semantically consistent responses and keeps extending its meaningless thinking until collapse. Detailed results can be found in Sec. A.7.1. Discussions on generalization over multi-purpose benchmarks can be found in Sec. A.7.2.
Table 2: Performance on ComplexBench (Qwen2.5-7B-Instruct). Best/2nd best are marked bold/underlined. OD, SC, CNFR, FC, and SR stand for the Oracle Decomposition wen2024benchmarking , Self-Consistency wang2022selfconsist , Conifer sun2024conifer , FollowComplex he2024complex , and Self-Refine madaan2023self .
@@
We implemented SOTAs on the ComplexBench (see Table 2): oracle decomposition wen2024benchmarking (ground-truth sub-instructions), self-consistency wang2022selfconsist (majority voting@10), Conifer sun2024conifer (SFT), FollowComplex he2024complex (SFT), and Self-Refine madaan2023self . Our method demonstrate its superiority on the most complicated Chain, Selection categories, suggesting that the reasoning indeed assists analysis of LLMs to carry out the truly relevant, valid request with constraints.
@@ A.2.1 Benchmarks
IFEval [7] is one commonly used benchmark to evaluate the abilities of LLMs in following natural language instructions. It focuses on a wide range of "verifiable instructions" that can be efficiently and accurately validated via python codes. It identifies 25 types of atomic constraints including keywords, languages, length constraints, detectable formats, combinations, case changes, starting and ending phrases, and punctuations. All the verifications are performed via codes to avoid potential bias caused by LLM judges. Therefore, their constraints are limited in types where no semantic ones are covered.
@@ A.2.2 Evaluation Metrics
ComplexBench [11] provides the scoring questions for each complex isntruction to check whether the constraints are satisfied. Each scoring question can only be answered with boolean responses as "YES" or "NO" and then the scores of each decomposed question are aggregated for the metric of decomposed requirements following ratio (DRFR). Note that the dependency of scoring questions is considered in aggregation. If any preceding constraint is not satisfied, the following constraints are simply judged as failure.
@@
InfoBench [54] proposes the decomposed requirements following ratio (DRFR) metric for verifications. For each scoring question, the LLM-as-a-Judge strategy is used to give a "YES" or "NO" response. The final scores of the response are simply the averaged accumulation of scores over the total number of scoring questions. Such scoring technique enables a more fine-grained interpretation of the final results.
@@ A.3.1 Tagging for Selection of Seed Instructions
For the selection of seed instructions, we choose the WildChat [34] (650K) and the Alpaca [86] (52K) as the pool of instructions. Especially for the WildChat dataset, a broad spectrum of user-chatbot interactions are covered including ambiguous user requests, code-switching, topic-switching, and political discussions. Therefore, it is indispensable to filter out those instructions that are not appropriate for constructing complex instructions with additional constraints. In this case, we follow InsTag [87] to perform topic and capability tagging on the instructions. We refer to the "intention tags" of InsTag as the ability tags in the present study because those open-ended tags are closely associated with the abilities of LLMs to solve input queries. Specifically, we manually summarize and define the topic tags from the approximate 3.2K ability tags of the InsTag (see Fig. 6). Two typical examples of the instag tagging by TagLM are provided below respectively for Chinese and English instructions from the WildChat, respectively (see Fig. 7).
After tagging, we perform selection of seed instructions. We choose instructions that possess topic tags of "Problem Solving", "NLP and Understanding", "Info Processing", and "Creativity & Design" as candidates. Then, we perform random sampling on those instructions. During sampling, the ability tags are used to maximize the diversity of the chosen seed instructions (i.e., the number of instructions classified to each ability tag is controlled). Finally, we obtain 39K seed instructions in total.
@@
"User Query": "Please paraphase following words to be more formal and cohesive: Among a wide range of generation tasks, subject-driven generation is a popular topic, which is also known as customization or personalization. Such task refers to learning object-level representations from a reference set and regenerate them in new contexts, depending on different conditions or prompts.",
"InsTag Response": "{"Ability Tags":["paraphrase", "formal tone"], "Topic Tags":["NLP and Understanding"]}",
"User Query": "介绍一下Gradio和docker",
"InsTag Response": "{"Ability Tags":["programming language", "question answer", "problem-solving"], "Topic Tags":["Problem Solving", "Programming & Dev", "Knowledge Q&A"]}",
@@ A.3.2 Generation of Rules and Constraints with Verifications
"detectable_format:multiple_sections", "change_case:capital_word_frequency",
"detectable_format:number_highlighted_sections", "detectable_format:number_bullet_lists",
"detectable_content:postscript", "detectable_content:number_placeholders",
"length_constraints:number_words", "length_constraints:number_paragraphs",
"length_constraints:number_sentences","language:response_language",
"keywords:letter_frequency", "startend:end_checker"
@@
"keywords:existence", "keywords:frequency",
"keywords:forbidden_words", "length_constraints:nth_paragraph_first_word"
@@
Prompt to Generate Keywords for Atomic Constraint Templates in English You are provided with an <instruction>. Your object is to come up some keywords that may be used to answer the <instruction>. They are usually related to the task described in the <instruction>. you should output your thinking process and the keywords you come up with.
@@
the <instruction> as to explain GANs, hence, ’architecture’,’training’ and ’generator’ may be appropriate keywords to use in the answer.
@@
"architecture", "training", "generator"
@@
Prompt to Generate Keywords for Atomic Constraint Templates in Chinese You are provided with an <instruction>. Your object is to come up some keywords that may be used to answer the <instruction>. They are usually related to the task described in the <instruction>. you should output your thinking process and the keywords you come up with.
@@
- If there exists passages/snippets that are enclosed/wrapped by “‘ or ”’ or """ in the original sub-instructions, you MUST generate NEW passages/snippets that are enclosed/wrapped in a similar way.
@@
A Complex Instruction and its Code-Verifiable Constraints { "instructions": "Explain and justify the need to learn coding in school. State the benefits it has for the future. You will answer this question in relation to using Scratch. Your final product will be to create a game of your choice. You can explain the ATL skills it helps you develop (200 words) End your response with a postscript indicated by P.S.. Respond with at least 3 sentences. response should contain the keyword "digital". The words game cannot be in the response. Your answer must be in the form of exactly 4 bullet points with the format below:
@@
* This is bullet point 2.",
"instruction_id_list":"[[[[
"detectable_content:postscript",
"length_constraints:number_sentences",
"keywords:existence",
"keywords:forbidden_words",
"detectable_format:number_bullet_lists"
@@
A Complex Instruction and its Scoring Questions { "instructions": "You are tasked with conducting a detailed analysis of the economic impact of a new trade policy on a small town. Assume you are a researcher with a deep understanding of economic principles and historical context. Your analysis should be thorough, incorporating both quantitative data and qualitative insights. The town in question has recently implemented a policy to reduce tariffs on imported goods. Your task is to evaluate the potential benefits and drawbacks of this policy. Additionally, you need to provide recommendations for the town’s policymakers based on your findings. Ensure that your analysis is presented in a formal academic style, with clear and concise language.",
"scoring_questions":"[[[[
"Does the response demonstrate a deep understanding of economic principles and historical context?",
"Does the response include both quantitative data and qualitative insights?",
"Does the response present a clear and concise evaluation of the policy’s potential benefits and drawbacks?",
"Does the response provide actionable recommendations for the town’s policymakers?",
"Is the response presented in a formal academic style with clear and concise language?"
@@ A.3.3 Quality Check for Filtering of the Self-Evolved Instructions
Prompt to Detect Language Inconsistency 你是一个答案评判专家,负责判断<答案>是否完全符合<问题>中提及的要求。
@@
Prompt to Detect Answer Leakage 你是一个答案评判专家,负责判断<答案>是否完全符合<问题>中提及的要求。
@@
"应用趋势": {
"云原生与多云支持": {
"中文": "随着云原生技术的成熟和企业多云策略的普及,MongoDB在2023年进一步加强了对云原生和多云环境的支持,帮助企业更灵活地管理和迁移数据。",
"英文": "With the maturation of cloud-native technologies and the widespread adoption of multi-cloud strategies, MongoDB has further enhanced its support for cloud-native and multi-cloud environments in 2023, enabling businesses to manage and migrate data more flexibly."
"实时数据分析与处理": {
"中文": "2023年,MongoDB在实时数据分析和处理方面取得了显著进展,通过优化查询性能和引入新的分析工具,满足了企业对实时数据洞察的需求。",
"英文": "In 2023, MongoDB made significant progress in real-time data analysis and processing, improving query performance and introducing new analytical tools to meet the demand for real-time data insights in businesses."
"安全性与合规性": …
</示例问题>
<示例评判结果>
@@
Prompt to Detect Under- or Over-Length Responses 你是一个答案评判专家,负责判断<答案>是否完全符合<问题>中提及的要求。
@@
Prompt to Detect Hallucination 你是一个答案评判专家,负责判断<答案>是否完全符合<问题>中提及的要求。
@@
Prompt to Detect Poorly Defined Instructions 你是一个答案评判专家,负责判断<答案>是否完全符合<问题>中提及的要求。
@@
Prompt to Detect Problematic Instructions 你是一个答案评判专家,负责判断<答案>是否完全符合<问题>中提及的要求。
@@
</示例问题>
<示例答案>
"Das Lied ’Qing Tian’ von Jay Chou beginnt mit den Zeilen: ’Die Sonne lacht, der Himmel ist blau, die Wolken sind weiß wie Schnee. Du lächelst, und ich bin glücklich, aber ich wage es nicht, dich anzusehen. Die Zeit vergeht, die Jahreszeiten wechseln, aber ich kann nicht vergessen, wie du gelächelt hast. Die Erinnerung an diesen Tag, als der Himmel so klar war, wird für immer in meinem Herzen bleiben.’ Dies sind jedoch nicht die genauen Worte, da das Lied ursprünglich auf Chinesisch ist und diese Zeilen eine Übersetzung der Bedeutung sind.****** Das Lied ’Qing Tian’ von Jay Chou enthält folgende Zeilen: ’Die Sonne strahlt hell, der Himmel ist so blau, die Wolken sind wie Watte. Dein Lächeln ist so süß, es macht mich glücklich, aber ich traue mich nicht, dich direkt anzusehen. Die Zeit vergeht, die Jahreszeiten wechseln, aber die Erinnerung an dein Lächeln bleibt. Dieser klare Tag, an dem der Himmel so strahlend war, wird in meinem Herzen bleiben.’ Diese Zeilen sind eine Interpretation der Bedeutung des Liedes, da es ursprünglich auf Chinesisch geschrieben wurde."
</示例答案>
<示例评判结果>
@@
Prompt to Detect NSFW Contents 你是一个答案评判专家,负责判断<答案>是否完全符合<问题>中提及的要求。
@@
In your entire response, the letter a should appear at least 5 times. give your answer wrapped in double angular brackets, such as <<your answer>>. organize your entire response in at least 3 sentences. Do not include the keywords: Ahri. Please limit the number of words with all capital letters to less than 2.
</示例问题>
<示例答案>
<<In a realm far from the eyes of many, Syndra, a powerful and enigmatic figure from the League of Legends, found herself in an unusual predicament. While she was lost in the depths of slumber, an unseen hand placed a chastity cage around her unexpected anatomy, a secret known to very few. The dawn broke, and Syndra awoke to find her situation, a challenge that would test her control and patience, as she navigated her day with a newfound restraint.>>
</示例答案>
<示例评判结果>

### v2

@@ 4.1 Experimental Setup
Details on statistics about our self-evolving dataset can be found in Sec. A.3. Additionally, we incorporated the DeepScaleR Luo et al. (2025) maths dataset to facilitate reasoning development (see Sec. A.4). The evaluation metrics for each benchmark are reported in Sec. A.2.2.
@@
We compared with: 1) I/O: direct input with $x_{\text{CoT}}=\emptyset$; 2) CoT: reasoning prompting Wei et al. (2022) to first deliver the thought and then the answer; 3) SDC: self-DeepClaude ErlichLiu (2025); getAsterisk (2025) technique that first prompts for the thought and then packs the original input with the thought as a new context for the answer (see Sec. A.5.3); 4) SFT: supervised fine-tuning for learning the aligned responses.
@@
We use OpenRLHF Hu et al. (2024) for both cold-start (Qwen2.5-1.5B/7B Yang et al. (2024c), LLaMA3.1-8B Grattafiori et al. (2024), and Ministral-8B Jiang et al. (2023b)) and warm-start (DeepSeek-Qwen1.5B/7B Guo et al. (2025) and DeepScaleR-1.5B Luo et al. (2025)) experiments. Detailed settings can be found in Sec. A.5. To ensure fairness and comparability, we employ Qwen2.5-72B-Instruct as the judge for all LLM-based evaluation.
@@ 4.2 Main Results
In line with Fig. 3, small models (1.5B) achieve much more gains than larger ones, showcasing the potentials of small LLMs via test-time scaling. The DeepSeek-distilled LLMs possess a good starting point for reasoning organization from their warm-start imitation across a broad range of tasks and topics (Fig. 36). With respect to model family, we unfortunately find that the capacity of Ministral and LLaMA is inferior to that of Qwen. The Ministral-8B exhibits limited advantages over its vanilla counterpart while the LLaMA3.1-8B experienced a model collapse during training. As shown in Fig. 35(a) and (d), a rapid shrinkage of response and a frequent surge of KL penalty imply a great deviation of LLaMA from its initial state. The reason behind might be ascribed to the pre-trained knowledge of base models Gandhi et al. (2025). LLaMA tends to generate endless thinking without conforming to the required format. It then struggles to output semantically consistent responses and keeps extending its meaningless thinking until collapse. Detailed results can be found in Sec. A.7.1. Discussions on generalization over multi-purpose benchmarks can be found in Sec. A.7.2.
Table 2: Performance on ComplexBench (Qwen2.5-7B-Instruct). Best/2nd best are marked bold/underlined. OD, SC, CNFR, FC, and SR stand for the Oracle Decomposition Wen et al. (2024), Self-Consistency Wang et al. (2022b), Conifer Sun et al. (2024a), FollowComplex He et al. (2024b), and Self-Refine Madaan et al. (2023).
@@
We implemented SOTAs on the ComplexBench (see Table 2): oracle decomposition Wen et al. (2024) (ground-truth sub-instructions), self-consistency Wang et al. (2022b) (majority voting@10), Conifer Sun et al. (2024a) (SFT), FollowComplex He et al. (2024b) (SFT), and Self-Refine Madaan et al. (2023). Our method demonstrate its superiority on the most complicated Chain, Selection categories, suggesting that the reasoning indeed assists analysis of LLMs to carry out the truly relevant, valid request with constraints.
@@ 5 Conclusion
Contributions
Authors
Yulei Qin1 Gang Li1 Zongyi Li1 Zihan Xu1  Yuchen Shi1 Zhekai Lin1,2 Xiao Cui3 Ke Li1 Xing Sun1
Affiliations
1Tencent YouTu Lab 2Xiamen University 3The Chinese University of Hong Kong
@@ A.2.1 Benchmarks
IFEval [7] is one commonly used benchmark to evaluate the abilities of LLMs in following natural language instructions. It focuses on a wide range of ”verifiable instructions” that can be efficiently and accurately validated via python codes. It identifies 25 types of atomic constraints including keywords, languages, length constraints, detectable formats, combinations, case changes, starting and ending phrases, and punctuations. All the verifications are performed via codes to avoid potential bias caused by LLM judges. Therefore, their constraints are limited in types where no semantic ones are covered.
@@ A.2.2 Evaluation Metrics
ComplexBench [11] provides the scoring questions for each complex isntruction to check whether the constraints are satisfied. Each scoring question can only be answered with boolean responses as ”YES” or ”NO” and then the scores of each decomposed question are aggregated for the metric of decomposed requirements following ratio (DRFR). Note that the dependency of scoring questions is considered in aggregation. If any preceding constraint is not satisfied, the following constraints are simply judged as failure.
@@
InfoBench [54] proposes the decomposed requirements following ratio (DRFR) metric for verifications. For each scoring question, the LLM-as-a-Judge strategy is used to give a ”YES” or ”NO” response. The final scores of the response are simply the averaged accumulation of scores over the total number of scoring questions. Such scoring technique enables a more fine-grained interpretation of the final results.
@@ A.3.1 Tagging for Selection of Seed Instructions
For the selection of seed instructions, we choose the WildChat [34] (650K) and the Alpaca [86] (52K) as the pool of instructions. Especially for the WildChat dataset, a broad spectrum of user-chatbot interactions are covered including ambiguous user requests, code-switching, topic-switching, and political discussions. Therefore, it is indispensable to filter out those instructions that are not appropriate for constructing complex instructions with additional constraints. In this case, we follow InsTag [87] to perform topic and capability tagging on the instructions. We refer to the ”intention tags” of InsTag as the ability tags in the present study because those open-ended tags are closely associated with the abilities of LLMs to solve input queries. Specifically, we manually summarize and define the topic tags from the approximate 3.2K ability tags of the InsTag (see Fig. 6). Two typical examples of the instag tagging by TagLM are provided below respectively for Chinese and English instructions from the WildChat, respectively (see Fig. 7).
After tagging, we perform selection of seed instructions. We choose instructions that possess topic tags of ”Problem Solving”, ”NLP and Understanding”, ”Info Processing”, and ”Creativity & Design” as candidates. Then, we perform random sampling on those instructions. During sampling, the ability tags are used to maximize the diversity of the chosen seed instructions (i.e., the number of instructions classified to each ability tag is controlled). Finally, we obtain 39K seed instructions in total.
@@
”User Query”: ”Please paraphase following words to be more formal and cohesive: Among a wide range of generation tasks, subject-driven generation is a popular topic, which is also known as customization or personalization. Such task refers to learning object-level representations from a reference set and regenerate them in new contexts, depending on different conditions or prompts.”,
”InsTag Response”: ”{”Ability Tags”:[”paraphrase”, ”formal tone”], ”Topic Tags”:[”NLP and Understanding”]}”,
”User Query”: ”介绍一下Gradio和docker”,
”InsTag Response”: ”{”Ability Tags”:[”programming language”, ”question answer”, ”problem-solving”], ”Topic Tags”:[”Problem Solving”, ”Programming & Dev”, ”Knowledge Q&A”]}”,
@@ A.3.2 Generation of Rules and Constraints with Verifications
”detectable_format:multiple_sections”, ”change_case:capital_word_frequency”,
”detectable_format:number_highlighted_sections”, ”detectable_format:number_bullet_lists”,
”detectable_content:postscript”, ”detectable_content:number_placeholders”,
”length_constraints:number_words”, ”length_constraints:number_paragraphs”,
”length_constraints:number_sentences”,”language:response_language”,
”keywords:letter_frequency”, ”startend:end_checker”
@@
”keywords:existence”, ”keywords:frequency”,
”keywords:forbidden_words”, ”length_constraints:nth_paragraph_first_word”
@@
Prompt to Generate Keywords for Atomic Constraint Templates in English You are provided with an ¡instruction¿. Your object is to come up some keywords that may be used to answer the ¡instruction¿. They are usually related to the task described in the ¡instruction¿. you should output your thinking process and the keywords you come up with.
@@
the ¡instruction¿ as to explain GANs, hence, ’architecture’,’training’ and ’generator’ may be appropriate keywords to use in the answer.
@@
”architecture”, ”training”, ”generator”
@@
Prompt to Generate Keywords for Atomic Constraint Templates in Chinese You are provided with an ¡instruction¿. Your object is to come up some keywords that may be used to answer the ¡instruction¿. They are usually related to the task described in the ¡instruction¿. you should output your thinking process and the keywords you come up with.
@@
- If there exists passages/snippets that are enclosed/wrapped by “‘ or ”’ or ””” in the original sub-instructions, you MUST generate NEW passages/snippets that are enclosed/wrapped in a similar way.
@@
A Complex Instruction and its Code-Verifiable Constraints { ”instructions”: ”Explain and justify the need to learn coding in school. State the benefits it has for the future. You will answer this question in relation to using Scratch. Your final product will be to create a game of your choice. You can explain the ATL skills it helps you develop (200 words) End your response with a postscript indicated by P.S.. Respond with at least 3 sentences. response should contain the keyword ”digital”. The words game cannot be in the response. Your answer must be in the form of exactly 4 bullet points with the format below:
@@
* This is bullet point 2.”,
”instruction_id_list”:”[[[[
”detectable_content:postscript”,
”length_constraints:number_sentences”,
”keywords:existence”,
”keywords:forbidden_words”,
”detectable_format:number_bullet_lists”
@@
A Complex Instruction and its Scoring Questions { ”instructions”: ”You are tasked with conducting a detailed analysis of the economic impact of a new trade policy on a small town. Assume you are a researcher with a deep understanding of economic principles and historical context. Your analysis should be thorough, incorporating both quantitative data and qualitative insights. The town in question has recently implemented a policy to reduce tariffs on imported goods. Your task is to evaluate the potential benefits and drawbacks of this policy. Additionally, you need to provide recommendations for the town’s policymakers based on your findings. Ensure that your analysis is presented in a formal academic style, with clear and concise language.”,
”scoring_questions”:”[[[[
”Does the response demonstrate a deep understanding of economic principles and historical context?”,
”Does the response include both quantitative data and qualitative insights?”,
”Does the response present a clear and concise evaluation of the policy’s potential benefits and drawbacks?”,
”Does the response provide actionable recommendations for the town’s policymakers?”,
”Is the response presented in a formal academic style with clear and concise language?”
@@ A.3.3 Quality Check for Filtering of the Self-Evolved Instructions
Prompt to Detect Language Inconsistency 你是一个答案评判专家,负责判断¡答案¿是否完全符合¡问题¿中提及的要求。
@@
Prompt to Detect Answer Leakage 你是一个答案评判专家,负责判断¡答案¿是否完全符合¡问题¿中提及的要求。
@@
”应用趋势”: {
”云原生与多云支持”: {
”中文”: ”随着云原生技术的成熟和企业多云策略的普及,MongoDB在2023年进一步加强了对云原生和多云环境的支持,帮助企业更灵活地管理和迁移数据。”,
”英文”: ”With the maturation of cloud-native technologies and the widespread adoption of multi-cloud strategies, MongoDB has further enhanced its support for cloud-native and multi-cloud environments in 2023, enabling businesses to manage and migrate data more flexibly.”
”实时数据分析与处理”: {
”中文”: ”2023年,MongoDB在实时数据分析和处理方面取得了显著进展,通过优化查询性能和引入新的分析工具,满足了企业对实时数据洞察的需求。”,
”英文”: ”In 2023, MongoDB made significant progress in real-time data analysis and processing, improving query performance and introducing new analytical tools to meet the demand for real-time data insights in businesses.”
”安全性与合规性”: …
¡/示例问题¿
¡示例评判结果¿
@@
Prompt to Detect Under- or Over-Length Responses 你是一个答案评判专家,负责判断¡答案¿是否完全符合¡问题¿中提及的要求。
@@
Prompt to Detect Hallucination 你是一个答案评判专家,负责判断¡答案¿是否完全符合¡问题¿中提及的要求。
@@
Prompt to Detect Poorly Defined Instructions 你是一个答案评判专家,负责判断¡答案¿是否完全符合¡问题¿中提及的要求。
@@
Prompt to Detect Problematic Instructions 你是一个答案评判专家,负责判断¡答案¿是否完全符合¡问题¿中提及的要求。
@@
¡/示例问题¿
¡示例答案¿
”Das Lied ’Qing Tian’ von Jay Chou beginnt mit den Zeilen: ’Die Sonne lacht, der Himmel ist blau, die Wolken sind weiß wie Schnee. Du lächelst, und ich bin glücklich, aber ich wage es nicht, dich anzusehen. Die Zeit vergeht, die Jahreszeiten wechseln, aber ich kann nicht vergessen, wie du gelächelt hast. Die Erinnerung an diesen Tag, als der Himmel so klar war, wird für immer in meinem Herzen bleiben.’ Dies sind jedoch nicht die genauen Worte, da das Lied ursprünglich auf Chinesisch ist und diese Zeilen eine Übersetzung der Bedeutung sind.****** Das Lied ’Qing Tian’ von Jay Chou enthält folgende Zeilen: ’Die Sonne strahlt hell, der Himmel ist so blau, die Wolken sind wie Watte. Dein Lächeln ist so süß, es macht mich glücklich, aber ich traue mich nicht, dich direkt anzusehen. Die Zeit vergeht, die Jahreszeiten wechseln, aber die Erinnerung an dein Lächeln bleibt. Dieser klare Tag, an dem der Himmel so strahlend war, wird in meinem Herzen bleiben.’ Diese Zeilen sind eine Interpretation der Bedeutung des Liedes, da es ursprünglich auf Chinesisch geschrieben wurde.”
¡/示例答案¿
¡示例评判结果¿
@@
Prompt to Detect NSFW Contents 你是一个答案评判专家,负责判断¡答案¿是否完全符合¡问题¿中提及的要求。
@@
In your entire response, the letter a should appear at least 5 times. give your answer wrapped in double angular brackets, such as ¡¡your answer¿¿. organize your entire response in at least 3 sentences. Do not include the keywords: Ahri. Please limit the number of words with all capital letters to less than 2.
¡/示例问题¿
¡示例答案¿
¡¡In a realm far from the eyes of many, Syndra, a powerful and enigmatic figure from the League of Legends, found herself in an unusual predicament. While she was lost in the depths of slumber, an unseen hand placed a chastity cage around her unexpected anatomy, a secret known to very few. The dawn broke, and Syndra awoke to find her situation, a challenge that would test her control and patience, as she navigated her day with a newfound restraint.¿¿
¡/示例答案¿
¡示例评判结果¿